7. Find the sum of 12 terms of an A.P. whose nth term is given by an = 3n + 4
(a) 262
(b) 272
(c) 282
(d) 292

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Answer: a Explaination:Reason: Here an = 3n + 4 ∴ a1 = 7, a2 – 10, a3 = 13 ∴ a= 7, d = 10 – 7 = 3 ∴ S12 = 122[2 × 7 + (12 – 1) ×3] = 6[14 + 33] = 6 × 47 = 282